News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 19 September 2003 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Real-time procedures were completed on 12 Sept to swap back from SSR-B to SSR-A after a DSN outage led to two missed passes and an autonomous switch to SSR-B. JPL resolved the problem and all data were recovered during the following pass. A real-time procedure was completed on Feb17 to update the on-board ephemeris.

The 4 Years of Chandra Observations, A Tribute to Riccardo Giacconi Symposium was held in Huntsville Sept 16-18. In addition to the excellent science program, highlights included Dr. Giacconi's address and a visit by STS-93 crew Cady Coleman and Jeff Ashby.

Two press releases were issued in conjunction with the Symposium. The releases titled 'Luna Prospecting with Chandra' and '"Iron-Clad" evidence For Spinning Black Hole' can be found at: http://chandra.harvard.edu/press/03_releases/press_091603.html. http://chandra.harvard.edu/press/03_releases/press_091703.html.

The observing schedule for next week is shown below and includes a calibration observation of Capella coordinated with XMM-Newton.

------------------------------------------
44I BOO                HRC-S/LETG   Sep 21
NGC 5204 X-1           ACIS-S       Sep 23
HR 6200                HRC-I
Radiation Belts
NGC 7619               ACIS-S       Sep 24
CL1641+4001            ACIS-I
4U 1323-619            ACIS-S/HETG  Sep 25
4U1755-33              ACIS-S
V4743 Sgr              HRC-S/LETG
Radiation Belts
Crab Pulsar            HRC-S/LETG   Sep 26
Eta Carinae            ACIS-S/HETG
Capella                ACIS-S/HETG  Sep 27
Capella                HRC-S/LETG   Sep 28
B2 1615+35             ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports